A woman owns stock valued at $2500. One day the stock drops by 6%, and then gains the same percentage back the next day. How much money is the stock valued at the end of the second day?
$____

Answer by ikleyn(52752)

.
At the first day evening the stock valued 2500 - 2500*0.06 = 2500 - 150 = 2350 dollars.

At the second day evening the stock valued 2350 + 0.06*2350 = 2350 + 141 = 2491.

You calculate the difference.


---------------
There is a bunch of lessons on percentage problems
    - Percentage problems
    - Percentage word problems (Type 1 problems, Finding the Part)
    - Percentage word problems (Type 2 problems, Finding the Rate)
    - Percentage word problems (Type 3 problems, Finding the Base)
    - More complicated percentage problems
    - Problems on percentage that lead to unexpected results
in this site.